OpenStudy (anonymous):

How do I simplify 15 over 2 ?

OpenStudy (anonymous):

\[\frac{ 15 }{ 2}\]

OpenStudy (anonymous):

How many times does 15 go into 2?

OpenStudy (anonymous):

Sorry, *2 into 15

OpenStudy (anonymous):

6 times.

OpenStudy (anonymous):

No, try again.

OpenStudy (anonymous):

2, 4, 6, 8, 10, 12, 14

OpenStudy (anonymous):

Yeah but if you go in to it 7x it will be 16

OpenStudy (anonymous):

8x2=15 7x2=?

OpenStudy (anonymous):

Sorry, 8x2=16

OpenStudy (anonymous):

7x2 is 14

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

So what would the whole number be?

OpenStudy (anonymous):

Here, I'll help you out.

OpenStudy (anonymous):

So because 2 goes into 15 7 times the whole number would 7 \[7 \frac{ ? }{ 2 }\]

OpenStudy (anonymous):

So because you know 2x7=14 what is left over of that to be the numerator?

OpenStudy (anonymous):

so it will be ?\[7\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No problem :)